OpenStudy (anmsteelers):

The midpoint of CD is E(-1,0). One endpoint is C(5,2). What are the coordinates of the other endpoint?

OpenStudy (chainn):

Let the x-coordinate of D be denoted by x while the y-coordinate of D be donated by y. (x-1)/2=5 [after solving you should get x=11) (y-0)/2=2 [after solving you should get y=4] Hence coordinates of D is (11, 4)

OpenStudy (anmsteelers):

Thank you!